# Enhancing LLM Reliability via Explicit Knowledge Boundary Modeling

### v2

@@ 5.1.2 Dataset
We evaluate our methods on the Dialogue State Tracking (DST) task in task-oriented dialogue systems, which involves extracting slot-value pairs from multi-turn dialogues based on a predefined ontology Xu et al. (2024a). Common evaluation metrics for DST include Joint Goal Accuracy (JGA) and Slot-F1. For JGA, a sample is deemed correct only if all predictions are accurate without omissions, while Slot-F1 offers a slot-level assessment.
To ensure a robust evaluation, we select three datasets: MultiWOZ-2.4 Ye et al. (2021), BiTOD Lin et al. , and SGD Rastogi et al. (2020). For BiTOD we only utilize the English version. For SGD, we randomly select 10,000 test samples.
5.1.3 Training Details
We modify the DeepSpeed-Chat Yao et al. (2023) framework to accommodate our experimental setup. We adopt most of the default training parameters provided by DeepSpeed-Chat except that the learning rate for DPO training is set to 1e-7 and the maximum sequence length is set to 8096 on the SGD dataset and 2048 on the others. By default, we utilize 1,000samples for Reliability-SFT Training and 2,000samples for DPO Training. All models are trained for 1 epoch using NVIDIA A800 GPUs.
@@ 5.2 Reliability Evaluation
The results of our experiments on model reliability are summarized in Table 1. We compare our reliability-training pipeline with baseline algorithms, emphasizing the model’s self-awareness ability. For RQ1, our method significantly outperforms traditional prompt-based and uncertainty-based methods across various datasets, markedly enhancing self-awareness.
Our approach consistently achieves higher Quality-F1 scores, indicating improved effectiveness in categorizing predictions as "sure" or "unsure" based on confidence. Notably, a high Precisionsure indicates reliable "sure" predictions, fostering user trust with minimal verification. Simultaneously, a high Recalltotal ensures that the "unsure" category encompasses a broad range of potentially correct outputs, forming a solid basis for further refinement. Collectively, these metrics contribute to a higher Optimal-F1 score, representing the theoretical maximum post-refinement performance. In other words, our system offers improved performance by effectively balancing precision and recall, ensuring both immediate reliability and the potential for improved post-refinement accuracy.
In our approach, methods incorporating DPO generally outperform SFT-only methods. Specifically, models trained with DPO Joint Training achieve higher Optimal-F1 scores, while those undergoing Post Training yield superior Quality-F1 scores. This suggests that Joint Training models are more conservative, relying on subsequent refinement, whereas Post Training models exhibit greater confidence, yielding predictions of generally higher quality. This discrepancy arises from the differing sampling distributions of DPO data; Joint Training samples from a pre-SFT distribution, where performance is relatively poor, leading to a dataset with a higher proportion of "unsure" predictions, resulting in more conservative model behavior. We will further explore the differences between these two approaches in subsequent sections.
Notably, some baselines, such as Self-Consistency, often achieve higher , , and Optimal-F1 scores compared to others, suggesting initial superiority. However, this indicates an undesirable scenario: an overly conservative model that frequently labels predictions as "unsure." This approach leads to a suboptimal balance between prediction types, evidenced by its significantly lower Quality-F1 score. In extreme cases, a model could assign "unsure" to all predictions, thereby maximizing Optimal-F1.However, this would merely shift the burden to the refinement stage, reduce immediate usability and fails to balance reliability and helpfulness.
@@ 5.3 Refinement for Unsure Prediction
We trained an automatic refinement model for each dataset as described in Section 4.3. We refine the predictions labeled as "unsure" from the initial stage and merge the refined outputs with the "sure" predictions to generate the final results. The overall evaluation results are presented in Table 2. For RQ2, It can be observed that our method, after refinement, achieves significantly better performance with substantial improvements in Slot-F1 and JGA.
Method Type Method MultiWOZ-2.4 BiTOD SGD LLAMA3-8B Prompt Direct 78.86 36.01 82.40 36.46 32.41 13.58 Verbose 76.80 36.69 76.93 27.99 23.11 9.00 SR 78.68 29.72 76.50 30.74 29.90 11.25 Uncertainty Prob 75.89 24.19 73.88 24.86 28.48 9.63 SC 70.98 20.70 69.98 24.86 25.60 4.39 P(True) 73.88 19.18 67.07 19.03 26.29 7.34 Reliability (Ours) SFT 85.95 53.76 85.21 60.09 74.04 27.62 + DPO Joint 86.47 54.31 84.53 61.37 75.47 29.18 + DPO Post 86.55 56.33 85.87 64.95 73.96 28.13 Qwen2.5-7B Prompt Direct 78.66 36.35 83.23 37.39 42.02 16.93 Verbose 67.59 26.45 74.37 27.38 27.79 9.19 SR 68.88 19.09 67.18 19.15 34.77 8.78 Uncertainty Prob 74.18 23.56 73.97 25.64 35.91 10.81 SC 59.34 23.45 65.60 28.31 34.34 3.06 P(True) 67.52 17.05 66.25 17.99 33.66 8.07 Reliability (Ours) SFT 84.50 51.90 85.29 58.39 74.86 28.02 + DPO Joint 87.06 52.41 85.44 62.31 74.85 27.06 + DPO Post 87.57 53.19 85.86 63.61 74.95 28.51
It is important to note that, referring to Section 5.2, despite a higher theoretical performance limit (Optimal-F1), DPO Joint Training models exhibit suboptimal Slot-F1 and JGA after refinement to Post Training models. This discrepancy stems from the limitations of the refinement model. Since the refinement model is not ideal, excessive "unsure" predictions can lead to residual errors after refinement, incurring significant costs without proportional gains. This highlights the importance of a balanced "sure" and "unsure" classification. An ideal model should minimize unnecessary "unsure" predictions and refine only when truly warranted, maximizing both efficiency and overall performance.
@@ 5.4 Scalability Analysis
To evaluate the scalability of our methods, we conducted experiments on foundation models with varying task capabilities. We regard the original LLAMA3 8B model as a lower-performance model (denoted as Low) and applied supervised fine-tuning (SFT) on a subset of the dataset to obtain models with improved performance. We trained two foundation models using 1,000 samples (Medium) and 10,000 samples (High) and compared their performance against baseline algorithms, with results shown in Figure 3. For RQ3, our methods consistently outperform baselines across different foundation models after refinement. Detailed results are available in Appendix C, leading to consistent conclusions.
Additionally, as illustrated in Appendix Table 4, the advantage of our DPO Post Training method over Joint Training increases with the improvement of foundation model performance. This is due to the offline nature of our DPO algorithm; as foundational model performance improves, the limitations of the Joint Training data distribution relative to Post Training become more pronounced, impacting overall performance. Notably, in the SGD dataset, DPO Joint Training becomes less beneficial as the foundation model’s performance rises.
5.5 Cost Analysis
In this work, "Cost" refers to the additional overhead from the refinement process. We define the cost of refining an "unsure" prediction as a constant value of 1, equating Cost to the number of "unsure" predictions. For better comparison, we express Cost as the proportion of "unsure" predictions.
We compare the performance and Cost of various baselines, as shown in Figure 4. Results are presented only for the MultiWOZ-2.4 dataset, with complete findings available in Appendix D. Our methods demonstrate significant performance improvements while maintaining a relatively low Cost. Furthermore, as illustrated in Appendix Table 3, Cost decreases significantly with improved foundation model performance. This aligns with intuitive expectations: as the model strengthens, the proportion of uncertain predictions diminishes.
@@ 5.6 DPO Preference Strategies
We construct our DPO preference data using Weighted-F1 as the ranking metric, as detailed in Section 3.2.1. Weighted-F1 includes two adjustable parameters: influences Weighted Precision, while affects Weighted Recall. Using the LLAMA3 8B Reliability-SFT model, we generated DPO preference data with various values and conducted Post Training. Test results on the MultiWOZ-2.4 dataset are shown in Figures 5(a) and 5(b). The model achieves optimal performance with set to 0.25 and to 0.75, reaching near-best results at a relatively low cost.
@@
Analysis indicates that increasing makes Weighted Precision more sensitive to errors in "unsure" predictions, reducing tolerance and generating fewer "unsure" outputs. In contrast, increasing raises the reward for successful recall from "unsure" predictions, encouraging the model to produce more "unsure" predictions and adopt a more conservative stance. Practically, the optimal alpha settings depend on the scenario and refinement model performance. When the refinement model is underperforming, it may be unwise to incur significant costs for marginal performance gains. The flexible adjustment of both parameters allows for adaptation to diverse real-world situations.
5.7 Refinement Model comparison
@@
We employ various refinement models to enhance the predictions of the LLAMA3 Reliability-SFT+DPO Post model, comparing their performance as shown in Table 6. In addition to our trained refinement models, we utilize GPT-4o and the recently popular DeepSeek series DeepSeek-AI et al. (2025), including the distilled 8B and 70B versions, all exhibiting strong reasoning capabilities. The results indicate significant performance variations among the different refinement models, with the DeepSeek 8B model underperforming, while the 70B model closely matches GPT-4o. Furthermore, our task-specific fine-tuning with CoT substantially improves refinement accuracy.
@@ 6 Conclusion
7 Limitations
While our EKBM framework demonstrates promising results, several limitations warrant discussion. First, the refinement stage incurs computational overhead proportional to the volume of "unsure" predictions, which may hinder scalability in latency-sensitive applications. Second, the efficacy of refinement critically depends on the performance of the secondary model; suboptimal refinement can propagate errors, offsetting gains from uncertainty labeling. Third, the manual tuning of Weighted-F1 parameters (, ) introduces scenario-specific biases, limiting generalizability. Finally, our evaluation focuses on dialogue state tracking tasks—broader validation across diverse domains is necessary to assess universal applicability. Addressing these challenges could further enhance the framework’s practicality and robustness.
Appendix A Refinement Model
We trained an automated refinement model for each dataset. The refinement model conducts in-depth reasoning on unsure predictions based on task-specific criteria, operating within a defined action space. Our refinement models are built upon SFT of LLAMA3 8B utilizing data constructed with OpenAI’s GPT-4o.

### v3

@@ 4.1.2 Dataset
We evaluate our methods on the Dialogue State Tracking (DST) task in task-oriented dialogue systems, which involves extracting slot-value pairs from multi-turn dialogues based on a predefined ontology (Xu et al., 2024a). We focus on the intricate multi-slot problem, where the model is likely to exhibit varying confidence levels for different parts of responses to the same query. Consequently, we do not adopt traditional QA or mathematical datasets. To ensure a robust evaluation, we utilize three datasets: MultiWOZ-2.4 (Ye et al., 2021), BiTOD (Lin et al., 2021) and SGD (Rastogi et al., 2020). For BiTOD, we adopt the English version, while for SGD, we randomly select 10,000 instances from the testset to mitigate excessive computational overhead.
Common evaluation metrics for DST include Joint Goal Accuracy (JGA) and Slot-F1. For JGA, a sample is deemed correct only if all predictions are accurate without omissions, while Slot-F1 offers a slot-level assessment.
@@ 4.2 Reliability Evaluation
The experimental results on model reliability are summarized in Table 1. We evaluate our reliability-training pipeline against baseline algorithms, with a particular focus on the model’s self-awareness capabilities. For RQ1, our method significantly outperforms traditional prompt-based and uncertainty-based methods across various datasets, demonstrating a substantial improvement in self-awareness.
Our approach consistently achieves superior Quality-F1 scores, indicating improved effectiveness in distinguishing predictions as ”sure” or ”unsure” based on confidence levels. Notably, a high indicates reliable ”sure” predictions, fostering user trust with minimal need for verification. Simultaneously, a high ensures that the ”unsure” category encompasses a broad spectrum of potentially correct outputs, forming a solid basis for further refinement. Collectively, these metrics contribute to a higher Optimal-F1 score, representing the theoretical maximum post-refinement performance. In other words, our system optimally balances precision and recall, delivering both immediate reliability and greater potential for post-refinement accuracy.
In our approach, methods incorporating DPO generally outperform SFT-only methods. Specifically, models trained with DPO Joint Training achieve higher Optimal-F1 scores, while those undergoing Post Training yield superior Quality-F1 scores. This suggests that Joint Training models adopt a more conservative stance, deferring to subsequent refinement, whereas Post Training models exhibit greater confidence, producing predictions of generally higher quality. This discrepancy stems from the differing sampling distributions of DPO data: Joint Training samples from a pre-SFT distribution, where performance is relatively poor, leading to a dataset with a higher proportion of ”unsure” predictions and resulting in more cautious model behavior.
Notably, some baselines, such as Self-Consistency, often achieve higher , , and Optimal-F1 scores compared to others, suggesting initial superiority. However, this indicates an undesirable scenario: an overly conservative model that frequently labels predictions as ”unsure.” This approach leads to a suboptimal balance between prediction types, evidenced by its significantly lower Quality-F1 score. In extreme cases, a model could assign ”unsure” to all predictions, thereby maximizing Optimal-F1.However, this would merely shift the burden to the refinement stage, reduce immediate usability and fails to balance reliability and helpfulness.
@@ 4.3 Refinement for Unsure Prediction
We trained an automatic refinement model for each dataset as described in Section 3.3. We refine the predictions labeled as ”unsure” from the initial stage and merge the refined outputs with the ”sure” predictions to generate the final results. The overall evaluation results are presented in Table 2. For RQ2, It can be observed that our method, after refinement, achieves significantly better performance with substantial improvements in Slot-F1 and JGA.
Method Type | Method | MultiWOZ-2.4 | BiTOD | SGD | |||
LLAMA3-8B | |||||||
Prompt | Direct | 78.86 | 36.01 | 82.40 | 36.46 | 32.41 | 13.58 |
Verbose | 76.80 | 36.69 | 76.93 | 27.99 | 23.11 | 9.00 | |
SR | 78.68 | 29.72 | 76.50 | 30.74 | 29.90 | 11.25 | |
Uncertainty | Prob | 75.89 | 24.19 | 73.88 | 24.86 | 28.48 | 9.63 |
SC | 70.98 | 20.70 | 69.98 | 24.86 | 25.60 | 4.39 | |
P(True) | 73.88 | 19.18 | 67.07 | 19.03 | 26.29 | 7.34 | |
Reliability (Ours) | SFT | 85.95 | 53.76 | 85.21 | 60.09 | 74.04 | 27.62 |
+ DPO Joint | 86.47 | 54.31 | 84.53 | 61.37 | 75.47 | 29.18 | |
+ DPO Post | 86.55 | 56.33 | 85.87 | 64.95 | 73.96 | 28.13 | |
Qwen2.5-7B | |||||||
Prompt | Direct | 78.66 | 36.35 | 83.23 | 37.39 | 42.02 | 16.93 |
Verbose | 67.59 | 26.45 | 74.37 | 27.38 | 27.79 | 9.19 | |
SR | 68.88 | 19.09 | 67.18 | 19.15 | 34.77 | 8.78 | |
Uncertainty | Prob | 74.18 | 23.56 | 73.97 | 25.64 | 35.91 | 10.81 |
SC | 59.34 | 23.45 | 65.60 | 28.31 | 34.34 | 3.06 | |
P(True) | 67.52 | 17.05 | 66.25 | 17.99 | 33.66 | 8.07 | |
Reliability (Ours) | SFT | 84.50 | 51.90 | 85.29 | 58.39 | 74.86 | 28.02 |
+ DPO Joint | 87.06 | 52.41 | 85.44 | 62.31 | 74.85 | 27.06 | |
+ DPO Post | 87.57 | 53.19 | 85.86 | 63.61 | 74.95 | 28.51 |
It is important to note that, referring to §4.2, despite a higher theoretical performance limit (Optimal-F1), DPO Joint Training models exhibit suboptimal Slot-F1 and JGA after refinement to Post Training models. This discrepancy stems from the limitations of the refinement model. Since the refinement model is not ideal, excessive ”unsure” predictions can lead to residual errors after refinement, incurring significant costs without proportional gains. This highlights the importance of a balanced ”sure” and ”unsure” classification. An ideal model should minimize unnecessary ”unsure” predictions and refine only when truly warranted, maximizing both efficiency and overall performance.
@@ 4.4 Scalability Analysis
To evaluate the scalability of our methods, we conducted experiments on foundation models with varying task capabilities. We consider the original LLAMA3 8B model a lower-performance baseline (denoted as Low) and applied supervised fine-tuning (SFT) on each dataset to enhance its DST capabilities. We train two foundation models using 1,000 samples (Medium) and 10,000 samples (High), then continually conduct reliability training and evaluate the performance of the obtained reliable LLM against baseline algorithms, with results shown in Figure 3. For RQ3, our methods consistently outperform baselines across different foundation models after refinement, demonstrating strong scalability—particularly in real-world scenarios where models are often domain-specific tuned and possess high task capabilities. Detailed results in Appendix §E further support these findings.
Additionally, as illustrated in Appendix Table 5, the advantage of DPO Post Training over Joint Training becomes more pronounced as foundation model performance improves. This trend stems from the offline nature of our DPO algorithm—higher-performing foundation models exacerbate the limitations of the Joint Training data distribution compared to Post Training, thereby affecting overall performance. Notably, in the SGD dataset, the benefits of DPO Joint Training diminish as the foundation model’s performance increases, even underperforming the SFT-only approach.
@@ 4.5 DPO Preference Strategies
We construct our DPO preference data using Weighted-F1 as the ranking metric, as detailed in §2.2.1. Weighted-F1 incorporates two adjustable parameters: influences Weighted Precision, while affects Weighted Recall. Using the Reliability-SFT trained LLAMA3 8B model as backbone, we further generated DPO preference data with varying values and conducted consecutive Post Training. Test results on the MultiWOZ-2.4 dataset (Figure 4) show that the model achieves optimal performance with and , reaching near-best results at a relatively low cost.
@@
Analysis indicates that increasing heightens sensitivity to errors in ”unsure” predictions, reducing tolerance and yielding fewer ”unsure” outputs. Conversely, increasing amplifies rewards for successfully recalling ”unsure” predictions, encouraging the model to produce more ”unsure” predictions and adopt a more conservative stance. Practically, the optimal alpha settings depend on the scenario and refinement model performance. If the refinement model underperforms, excessive costs for marginal gains may not be justified. The flexible adjustment of both parameters allows for adaptation to diverse real-world scenarios.
4.6 Detailed Analysis
4.6.1 Prediction Types Analysis
We further analyzed the model’s behavior after Reliability Training to evaluate its ability to make reliable confidence judgments in its outputs. Specifically, we examined the distribution of output categories and compared them to baseline models. These categories include five cases: correct and incorrect predictions classified as ”sure” or ”unsure”, along with missing predictions. The results are presented in Figure 5.
Our findings highlight two key improvements: (1) Enhanced reliability—The proportion of incorrect predictions within the ”sure” category has been notably reduced, with most errors now concentrated in the ”unsure” category. This indicates a stronger self-assessment capability, allowing the model to better identify potentially erroneous predictions compared to baselines. Additionally, across all scenarios, the ”unsure” category contains fewer correct predictions than incorrect ones, demonstrating the model’s effectiveness in placing accurate outputs in the ”sure” category, a distinction baseline models struggle to achieve. (2) Improved recall—Our approach significantly reduces missing predictions, capturing previously omitted relevant information within the ”unsure” category. While some may still contain errors, this shift enhances overall system helpfulness and provides a stronger foundation for further refinement.
4.6.2 Cost Analysis
In this work, ”Cost” refers to the additional overhead from the refinement process. We define the cost of refining an ”unsure” prediction as a constant value of 1, equating Cost to the number of ”unsure” predictions. For better comparison, we express Cost as the proportion of ”unsure” predictions.
We compare the performance and Cost of various baselines, as shown in Figure 6. Results are presented only for the MultiWOZ-2.4 dataset, with complete findings available in Appendix §F. Our methods demonstrate significant performance improvements while maintaining a relatively low Cost. Furthermore, as illustrated in Appendix Table 7, Cost decreases significantly with improved foundation model performance. This aligns with intuitive expectations: as the model strengthens, the proportion of uncertain predictions diminishes.
4.6.3 Refinement Model comparison
@@
We employ various refinement models to enhance the predictions of the LLAMA3 Reliability-SFT+DPO Post model, comparing their performance as shown in Table 7. In addition to our trained refinement models, we utilize GPT-4o and the recently popular DeepSeek series (DeepSeek-AI et al., 2025), including the distilled 8B and 70B versions, all exhibiting strong reasoning capabilities. The results indicate significant performance variations among the different refinement models, with the DeepSeek 8B model underperforming, while the 70B model closely matches GPT-4o. Furthermore, our task-specific fine-tuning with CoT substantially improves refinement accuracy.
5 Related Work
The knowledge boundary of LLMs has emerged as a critical topic in recent research, highlighting their limitations in generating reliable outputs (Yin et al., 2024; Li et al., 2024b). Misalignment between model behavior and knowledge boundaries can result in factual hallucinations (Huang et al., 2023) and ambiguous responses (Liu et al., 2023).
Various methods have been proposed to assess these boundaries. Uncertainty-based approaches quantify prediction confidence through token probabilities (Manakul et al., 2023) and consistency (Chen & Mueller, 2024). Calibration strategies, including prompting (Tian et al., 2023) and fine-tuning (Tao et al., 2024) for improved confidence expression, align model confidence with prediction accuracy. Internal State Probing evaluates prediction factuality by analyzing model states, such as activations (Li et al., 2024a). While effective, these methods often rely on external post-hoc techniques, which are computationally expensive and lack integration with the model’s endogenous reasoning processes (Huang et al., 2024; Zhou et al., 2024).
To further align model behavior rather than merely estimate uncertainty, some studies guide models in self-reflection to minimize self-contradictions (Wei et al., 2022; Ji et al., 2023), albeit with increased computational overhead. More recently, RL-based approaches leverage uncertainty estimation to train reward models, enhancing LLM truthfulness by rejecting queries beyond their capabilities (Xu et al., 2024b; Chen et al., 2024; Xue et al., 2024). While effective in mitigating misinformation, these methods rely on the accuracy of uncertainty estimation algorithms and meanwhile may unintentionally reduce overall helpfulness.
@@ 6 Conclusion
Appendix A Training Details
We modify the DeepSpeed-Chat (Yao et al., 2023) framework to accommodate our experimental setup. We adopt most of the default training parameters provided by DeepSpeed-Chat except that the learning rate for DPO training is set to 1e-7 and the maximum sequence length is set to 8096 on the SGD dataset and 2048 on the others. All models are trained for 1 epoch using NVIDIA A800 GPUs.
Appendix B Training Data Construction
We provide detailed examples to illustrate our training data construction process, including model inputs and the expected outputs the model is intended to learn. The data construction for SFT is outlined in Table 3, while that for DPO is detailed in Table 4.
Input |
Dialog history: USER: Is there a modern European restaurant in the north? SYSTEM: I found city stop restaurant. They are located at Cambridge City Football Club Milton Road Chesterton. USER: That sounds lovely! SYSTEM: Would you like me to help you with reserving a table? USER: Maybe. Is that restaurant in the moderate price range? SYSTEM: No, it is in the expensive price range. Is this okay, or would you like to try another area? USER: How about a restaurant that serves Indian food? SYSTEM: There are several restaurants what part of town would you like it? USER: A restaurant in the north, please. Reliable dialog state: |
Gold Labels |
{’restaurant’: {’area’: ’north’, ’food’: ’indian’, ’pricerange’: ’moderate’}} |
LLM Original Predictions |
{’restaurant’: {’area’: ’north’, ’food’: ’european’, ’name’: ’Cambridge City Football Club Milton Road Chesterton’}} |
Constructed SFT Sample Output |
{’restaurant’: {’area’: {’value’: ’north’, ’confidence’: ’sure’}, ’food’: {’value’: ’European’, ’confidence’: ’unsure’}, ’pricerange’: {’value’: ’moderate’, ’confidence’: ’unsure’}, ’name’: {’value’: ’Cambridge City Football Club Milton Road Chesterton’, ’confidence’: ’unsure’}}} |
Explanation |
In the LLM’s original predictions, the value for area is correct, food is incorrect, name is extraneous, and pricerange is missing. Consequently, in the constructed SFT sample output, area is labeled as ”sure,” while the other three are marked as ”unsure,” with erroneous information left uncorrected. |
Input |
Dialog history: USER: Is there a modern European restaurant in the north? SYSTEM: I found city stop restaurant. They are located at Cambridge City Football Club Milton Road Chesterton. USER: That sounds lovely! SYSTEM: Would you like me to help you with reserving a table? USER: Maybe. Is that restaurant in the moderate price range? SYSTEM: No, it is in the expensive price range. Is this okay, or would you like to try another area? USER: How about a restaurant that serves Indian food? SYSTEM: There are several restaurants what part of town would you like it? USER: A restaurant in the north, please. Reliable dialog state: |
Gold Labels |
{’restaurant’: {’area’: ’north’, ’food’: ’indian’, ’pricerange’: ’moderate’}} |
Multi-Sampling With the Foundation LLM |
Constructed DPO Sample Output Pairs |
Chosen {’restaurant’: {’area’: {’value’: ’north’, ’confidence’: ’sure’}, ’food’: {’value’: ’indian’, ’confidence’: ’sure’}, ’pricerange’: {’value’: ’moderate’, ’confidence’: ’unsure’}}} Rejected {’restaurant’: {’area’: {’value’: ’north’, ’confidence’: ’sure’}, ’food’: {’value’: ’indian’, ’confidence’: ’sure’}}} |
Explanation |
Compared to the rejected sample, the chosen sample exhibits a higher recall for a correct pricerange within the ”unsure” category, aligning more closely with our alignment objectives. This is reflected in a higher Weighted-F1 score, justifying its selection over the rejected sample. |
Appendix C Refinement Model
We trained an automated refinement model for each dataset. The refinement model conducts in-depth step-by-step reasoning on unsure predictions based on task-specific criteria, operating within a defined action space. Our refinement models are built upon LLAMA3 8B utilizing SFT training with data constructed using OpenAI’s GPT-4o.